Calculate the percent ionization of 0.115 M lactic acid in a solution containing 7.5×10^-3 M sodium lactate.

Ka of lactic acid=1.4x10^-4
Answer #1

[Acid] = 0.115 M

[Salt] = 0.0075 M

Ka = 1.4 x 10-4

pKa = -log Ka = -log 1.4 x 10-4 = 3.854

pH = pKa + log [Salt]/[Acid]

pH = 3.854 + log 0.0075 M/0.115 M

pH = 2.668

pH = - log [H+]

[H+] = 10-pH

[H+] = 10-2.668

[H+] = 0.002146 M

% ionization = [0.002146 M/0.115 M] x 100 %

% ionization = 1.866 %